Find the area of a parallogram (rhombus) with one angle measure of 60 deg. and one diagnol measuring 15 and another measuring 10
Answer by macston(5194)

.
The diagonals of a rhombus form four right triangles with
half of each diagonal as legs, and the sides as hypotenuses.
The legs ((1/2 diagonals) constitute the base and height,
so area of each DISABLED_event_one= 1/2bh=1/2(1/2 diagonal)(1/2 other diagonal)
or(1/8)(diagonal times diagonal)and since there are four of
these triangles, the total area=(4)(1/8)(diagonal x diagonal)=
(diagonal 1 x diagonal 2)/2=(10)(15)/2=150/2=75 square units.
ANSWER: The area of the rhombus is 75 square units.